OpenStudy (anonymous):

Change r=9sin theta into rectangular form

OpenStudy (shubhamsrg):

what is rectangular form ?

OpenStudy (sirm3d):

multiply \(r=9\sin\theta\) by \(r\) then use \[x^2+y^2=r^2\] and \[y=r\sin \theta\]